OpenStudy (anonymous):

A graduated cylinder contains 32.0 cm3 of Hg. If the density of Hg (@ 25˚C) is 13.534 g/cm3, how many moles of Hg are in the container?

OpenStudy (anonymous):

@zepp

OpenStudy (zepp):

How many g of Hg there are in the cylinder?

OpenStudy (anonymous):

13.534g

OpenStudy (zepp):

That's the density, what about in the cylinder?

OpenStudy (anonymous):

PV = NRT

OpenStudy (anonymous):

use this

OpenStudy (lgbasallote):

i think it's more of that molar mass thingy..since there's density

OpenStudy (zepp):

Define P,V,N,R and T, @Yahoo! ?

OpenStudy (lgbasallote):

what was that formula again?

OpenStudy (anonymous):

oh..no pressure is not given hm!!!

OpenStudy (zepp):

\[\large \frac{13.534g}{1cm^3}=\frac{?g}{32cm^3}\]

OpenStudy (zepp):

There's 433.088 g of mercury in the cylinder.

OpenStudy (lgbasallote):

n = m/M i believe that's the formula yes? n is moles m is mass M is molar mass

OpenStudy (lgbasallote):

\[\text{mole} = \frac{\rho V}{M}\]

OpenStudy (lgbasallote):

or am i wrong?

OpenStudy (zepp):

From the periodic table, we know that 1 mole of Hg = 200.59 grams\[\large \frac{1mol}{200.59g}=\frac{?mol}{433.08g}\]

OpenStudy (zepp):

@lgbasallote That's the right formula.

OpenStudy (lgbasallote):

cool

OpenStudy (lgbasallote):

so it's a direct sub what's with this thingy then?

OpenStudy (zepp):

That's just a formula, I'm using the long method.

OpenStudy (zepp):

With ratio it's much easier to understand.

OpenStudy (zepp):

and they are quite the same thing so..

OpenStudy (anonymous):

so 2.15 amu?

OpenStudy (anonymous):

ok, im lost, so first you gotta convert from cm3 to g right?

OpenStudy (zepp):

mole*